# Gorenstein singularities with $\mathbb{G}_{m}$-action and moduli spaces of holomorphic differentials

### v1

@@ 2.3. Properties of the construction
$\mathcal{F}^{\lambda}H^{0}(C,m\omega_{C}^{\log})\coloneqq\bigcap_{i=1}^{n}\left\{s\in H^{0}(C,m\omega_{C}^{\log})\mid a_{i}\operatorname{ord}_{p_{i}}(s)\geq\lambda\right\}.$ (5)
@@
    $\bigoplus_{m\in\mathbb{N}}H^{0}(\mathcal{C},m\mathcal{L})$ $\cong\bigoplus_{m\in\mathbb{N}}\bigoplus_{\lambda\in\mathbb{Z}}\mathcal{F}^{\lambda}H^{0}(C,m\omega_{C}^{\log})t^{-\lambda},$ (6)
    $\bigoplus_{m\in\mathbb{N}}H^{0}(Y,m\omega_{Y}^{\log})$ $\cong\bigoplus_{m\in\mathbb{N}}\bigoplus_{\lambda\in\mathbb{Z}}\mathcal{F}^{\lambda}H^{0}(C,m\omega_{C}^{\log})/\mathcal{F}^{\lambda+1}H^{0}(C,m\omega_{C}^{\log}).$ (7)
@@
    The space $\mathbb{P}\mathcal{H}(\mu)_{(Y,q)}$ can be identified with the locus $\mathbb{P}{\rm Def}^{-}_{s}(Y,q)$ of smooth deformations with negative grading in the versal deformation space.
@@
    $Y_{\infty}=\frac{1}{\ell}\cdot\sum_{i=1}^{n}(m_{i}+1)y_{i}\sim\frac{1}{\ell}\cdot\omega_{Y}^{\log}.$ (9)
@@
${\tilde{\mathcal{C}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{f}$$\scriptstyle{\rho}$${\mathcal{C}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${C_{\mathbb{A}^{1}}}$
@@
For the first claim, the family $\mathcal{C}$ is a smoothing of $Y$. Since the ideal of the blowup is $\mathbb{G}_{m}$-equivariant with respect to the standard $\mathbb{G}_{m}$-action on $\mathbb{A}^{1}$, it induces a natural $\mathbb{G}_{m}$-action on $\mathcal{C}$ which satisfies the definition of test configurations in [BHJ17, Section 2]; see in particular [BHJ17, Example 2.5]. Additionally, the relative holomorphic one-form on $\tilde{\mathcal{C}}$ with zeros $\sum_{i=1}^{n}m_{i}\tilde{P}_{i}$ descends to $\mathcal{C}$ as a local generator for the relative dualizing bundle of $\mathcal{C}$ at $q$, which implies that $q$ is a Gorenstein singularity; see [CC24, Section 2] for a more detailed explanation.
@@
where $\operatorname{ord}_{Y_{i}}(\ell\tilde{C}_{0})=0$ since the proper transform $\tilde{C}_{0}$ does not contain $Y_{i}$. By the proof of [BHJ17, Lemma 5.17] and [ABB+23, Equation (4.3)], we conclude that
$\mathcal{F}^{\lambda}H^{0}(C,m\omega_{C}^{\log})=\bigcap_{i=1}^{n}\left\{s\in H^{0}(C,m\omega_{C}^{\log})\mid a_{i}\operatorname{ord}_{p_{i}}(s)\geq\lambda\right\}.$
@@
For the fourth claim, we refer to [Pin74] about the properties of the versal deformation spaces for singularities with $\mathbb{G}_{m}$-action; see also [Loo84b, Appendix]. In our case, recall that $\mathbb{P}{\rm Def}^{-}(Y,q)$ is the part of smooth deformations with negative grading. If $s(C,p_{1},\ldots,p_{n})=(Y,q)$, then by our construction, it is clear that $(C,p_{1},\ldots,p_{n})\in\mathbb{P}{\rm Def}^{-}(Y,q)$, where the $\mathbb{G}_{m}$-action is attractive to the singularity $q$ and hence the corresponding deformation direction has negative grading in Pinkham’s convention. Conversely, consider a smoothable Gorenstein curve singularity $(Y,q)$ with $\mathbb{G}_{m}$-action such that $Y$ is a projective curve with $n$ irreducible rational components $Y_{i}$ that meet at $q$ and such that $(Y,q)$ admits a $\mathbb{G}_{m}$-equivariant smoothing. A local generator $\eta$ of $\omega_{Y}$ at $q$ restricted to each branch $Y_{i}$ is a meromorphic differential $\eta_{i}=dt_{i}/t_{i}^{m_{i}+2}$ for some $m_{i}\geq 0$. Then after performing stable reduction, the associated (smooth) tail of $(Y,q)$ (in the sense of [Has00, Definition 3.1] and [FS13, Section 4.4]) is a curve $C$ with a canonical divisor $\sum_{i=1}^{n}m_{i}p_{i}$. This curve $C$ is attached at each $p_{i}$ to a rational component $E_{i}$ that maps to $Y_{i}$ in the process of stable reduction; see [Bat24, Section 1.4] for more details. In particular, $(C,p_{1},\ldots,p_{n})\in\mathbb{P}\mathcal{H}(\mu)$ for $\mu=(m_{1},\ldots,m_{n})$. Hence, the identification between $\mathbb{P}\mathcal{H}(\mu)_{(Y,q)}$ and $\mathbb{P}{\rm Def}^{-}_{s}(Y,q)$ is established.
@@
For the last claim, the divisor at infinity in (2) along each branch $Y_{i}$ is defined by $t$ and $s_{i}=1/t_{i}$. Due to the weighted blowup of the ideal $(x_{i},t^{a_{i}})$, where $a_{i}=\ell/(m_{i}+1)$, the restriction of $t$ to $Y_{i}$ picks up the point $y_{i}$ defined by $s_{i}=0$ with weight $1/a_{i}=(m_{i}+1)/\ell$, and the last equivalence relation of divisor classes follows from $\sum_{i=1}^{n}(m_{i}+1)y_{i}\sim\omega_{Y}^{\log}$ in the construction of the contraction map $f$. Alternatively, we can compute the divisor at infinity by using [Pin77, Section 5, Formula ($\ast$)]. Indeed, in the family $\tilde{\mathcal{C}}$, we can further blow up each node of type $x_{i}t_{i}=t^{a_{i}}$ in the central fiber for $i=1,\ldots,n$, replacing it by a chain of rational curves of length $a_{i}$. The resulting family thus provides a minimal good resolution of $\mathcal{C}$ in the setting of [Pin77, Section 2], where the related invariants can be read off easily. Finally, note that $\omega_{Y}^{\log}$ is the restriction of the (relatively) ample line bundle $\mathcal{L}$ to $Y$, where $f^{*}\mathcal{L}\cong\tilde{\mathcal{L}}$ induces the contraction as in (4). Therefore, $\omega_{Y}^{\log}$ is ample. The same conclusion holds for $C$ and its divisor at infinity in the negatively graded versal deformation space of $(Y,q)$; see [Loo84b, (A.5)]. ∎
@@ 2.4. The boundary of $\mathbb{P}{\rm Def}^{-}(Y,q)$
As seen in Example 2.1, besides smooth elliptic curves in the versal deformation space of the ordinary cusp, the rational nodal curve also appears. In general, given a Gorenstein curve singularity $(Y,q)$ with $\mathbb{G}_{m}$-action, a natural question is to determine which singular curves can appear in $\mathbb{P}{\rm Def}^{-}(Y,q)$. This question has been studied in various special situations, mostly for unibranched monomial singularities with symmetric semigroups, for local complete intersections, and for low genus cases; see [Stö93, CS13, CF23, CM25, Del90, LP19, Bat22]. Utilizing the ideas in Theorem 2.6, we can provide a unified description to answer this question.
Analogous to the usual stability condition for nodal curves, we will use the following notion for Gorenstein curves. Let $(C,p_{1},\ldots,p_{n})$ be a reduced Gorenstein curve with $n$ marked points $p_{1},\ldots,p_{n}$ contained in the smooth locus of $C$. We say that $(C,p_{1},\ldots,p_{n})$ is a stable pointed Gorenstein curve if $\omega_{C}^{\log}=\omega_{C}(p_{1}+\cdots+p_{n})$ is an ample line bundle class. Furthermore, given a partition $\mu=(m_{1},\ldots,m_{n})$ of $22g-$ 2, we say that the pointed curve $(C,p_{1},\ldots,p_{n})$ is of type $\mu$ if $\sum_{i=1}^{n}m_{i}p_{i}$ is a canonical divisor of $C$.
@@
Conversely, suppose $(C,p_{1},\ldots,p_{n})$ is a stable pointed Gorenstein curve with a differential $\omega$ as a section of the dualizing bundle of $C$, where the zeros of $\omega$ yield the canonical divisor $\sum_{i=1}^{n}m_{i}p_{i}$ of type $\mu$. Since each $p_{i}$ is a smooth point of $C$, the construction of the weighted blowup in Section 2.2 can be carried out in the same way. Additionally, $\omega_{C}^{\log}$ is ample by assumption, which implies that the twisted relative log dualizing bundle $\tilde{\mathcal{L}}$ can be defined as in (3) and remains relatively semiample. Since $\tilde{\mathcal{L}}$ is a subbundle of $\omega_{\tilde{\pi}}^{\log}\otimes\mathcal{O}_{\tilde{\mathcal{C}}}(\ell(\tilde{\mathcal{C}}_{0}+\sum_{i=1}^{n}E_{i}))$ and the base of the family is a DVR, it follows that $\tilde{\pi}_{*}\tilde{\mathcal{L}}^{k}$ is locally free for $k\gg 0$. Therefore, the contraction in (4) holds. We can also justify that the resulting singularity $q$ in the central fiber $Y$ after the contraction is Gorenstein. Indeed, the Residue Theorem holds for reduced singular curves as well; see [LP19, Lemma 1.1.2]. Consequently, the same argument as in [CC24, Section 2] implies that the multi-scale differential $\tilde{\eta}$ associated to the limit of $\omega$ in the central fiber of the weighted blowup can descend to $Y$ and locally generate its dualizing sheaf at the singularity $q$. The construction of the blowup and contraction as in Section 2.2 is $\mathbb{G}_{m}$-equivariant. Therefore, $(Y,q)$ is a Gorenstein curve with $\mathbb{G}_{m}$-action of type $\mu$, and its stability follows from Theorem 2.6 (v).
@@
Finally, the isomorphism of graded rings as in (7) follows from [Loo84b, (A.5)], where the class of the divisor at infinity is just $\omega^{\log}/\ell$. ∎
@@ 3.1.4. Unibranched planar singularities
where $p$ and $q$ are coprime. Its $\mathbb{P}{\rm Def}^{-}_{s}$ corresponds to the locus in the stratum $\mathbb{P}\mathcal{H}(2g-2)$, where $g=(p-1)(q-1)/2$, with the gap sequence $G=\{b_{1},\ldots,b_{g}\}$ consisting of positive integers that cannot be expressed as $pi+qj$ with $i,j\geq 0$. In this case,

### v2

@@ 2.3. Properties of the construction
(5) $\mathcal{F}^{\lambda}H^{0}(C,m\omega_{C}^{\log})\coloneqq\bigcap_{i=1}^{n}\left\{s\in H^{0}(C,m\omega_{C}^{\log})\mid a_{i}\operatorname{ord}_{p_{i}}(s)\geq\lambda\right\}.$
@@
    (6) $\bigoplus_{m\in\mathbb{N}}H^{0}(\mathcal{C},m\mathcal{L})$ $\cong\bigoplus_{m\in\mathbb{N}}\bigoplus_{\lambda\in\mathbb{Z}}\mathcal{F}^{\lambda}H^{0}(C,m\omega_{C}^{\log})t^{-\lambda},$
    (7) $\bigoplus_{m\in\mathbb{N}}H^{0}(Y,m\omega_{Y}^{\log})$ $\cong\bigoplus_{m\in\mathbb{N}}\bigoplus_{\lambda\in\mathbb{Z}}\mathcal{F}^{\lambda}H^{0}(C,m\omega_{C}^{\log})/\mathcal{F}^{\lambda+1}H^{0}(C,m\omega_{C}^{\log}).$
@@
    The space $\mathbb{P}\mathcal{H}(\mu)_{(Y,q)}$ can be identified with the locus $\mathbb{P}{\rm Def}^{-}_{s}(Y,q)$ of smooth deformations with negative grading in the miniversal deformation space.
@@
    (9) $Y_{\infty}=\frac{1}{\ell}\cdot\sum_{i=1}^{n}(m_{i}+1)y_{i}\sim\frac{1}{\ell}\cdot\omega_{Y}^{\log}.$
@@
[…]
@@
For the first claim, the family $\mathcal{C}$ is a smoothing of $Y$. Since the ideal of the blowup is $\mathbb{G}_{m}$-equivariant with respect to the standard $\mathbb{G}_{m}$-action on $\mathbb{A}^{1}$, it induces a natural $\mathbb{G}_{m}$-action on $\mathcal{C}$ which satisfies the definition of test configurations in [BHJ17, Section 2]; see in particular [BHJ17, Example 2.5]. Additionally, the relative holomorphic one-form on $\tilde{\mathcal{C}}$ with zeros $\sum_{i=1}^{n}m_{i}\tilde{P}_{i}$ descends to $\mathcal{C}$ as a local generator for the relative dualizing bundle of $\mathcal{C}$ at $q$, which implies that $q$ is a Gorenstein singularity; see [CC26, Section 2] for a more detailed explanation.
@@
where $\operatorname{ord}_{Y_{i}}(\ell\tilde{C}_{0})=0$ since the proper transform $\tilde{C}_{0}$ does not contain $Y_{i}$. By the proof of [BHJ17, Lemma 5.17] and [ABB+23, Equation (4.3)], we conclude that the subspace with weight $\geq\lambda$ in $H^{0}(C,m\omega_{C}^{\log})$ is
$\mathcal{F}^{\lambda}H^{0}(C,m\omega_{C}^{\log})=\bigcap_{i=1}^{n}\left\{s\in H^{0}(C,m\omega_{C}^{\log})\mid a_{i}\operatorname{ord}_{p_{i}}(s)\geq\lambda\right\}$
as defined in (5).
@@
For the fourth claim, we refer to [Pin74] about the properties of the miniversal deformation spaces for singularities with $\mathbb{G}_{m}$-action; see also [Loo84b, Appendix]. In our case, recall that $\mathbb{P}{\rm Def}^{-}_{s}(Y,q)$ is the part of smooth deformations with negative grading. If $s(C,p_{1},\ldots,p_{n})=(Y,q)$, then by our construction, it is clear that $(C,p_{1},\ldots,p_{n})\in\mathbb{P}{\rm Def}^{-}_{s}(Y,q)$, where the $\mathbb{G}_{m}$-action is attractive to the singularity $q$ and hence the corresponding deformation direction has negative grading in Pinkham’s convention. Conversely, consider a smoothable Gorenstein curve singularity $(Y,q)$ with $\mathbb{G}_{m}$-action such that $Y$ is a projective curve with $n$ irreducible rational components $Y_{i}$ that meet at $q$ and such that $(Y,q)$ admits a $\mathbb{G}_{m}$-equivariant smoothing. A local generator $\eta$ of $\omega_{Y}$ at $q$ restricted to each branch $Y_{i}$ is a meromorphic differential $\eta_{i}=dt_{i}/t_{i}^{m_{i}+2}$ for some $m_{i}\geq 0$. Then after performing stable reduction, the associated (smooth) tail of $(Y,q)$ (in the sense of [Has00, Definition 3.1] and [FS13, Section 4.4]) is a curve $C$ with a canonical divisor $\sum_{i=1}^{n}m_{i}p_{i}$. This curve $C$ is attached at each $p_{i}$ to a rational component $E_{i}$ that maps to $Y_{i}$ in the process of stable reduction; see [Bat24, Section 1.4] for more details. In particular, $(C,p_{1},\ldots,p_{n})\in\mathbb{P}\mathcal{H}(\mu)$ for $\mu=(m_{1},\ldots,m_{n})$. Hence, the identification between $\mathbb{P}\mathcal{H}(\mu)_{(Y,q)}$ and $\mathbb{P}{\rm Def}^{-}_{s}(Y,q)$ is established.
@@
For the last claim, the divisor at infinity in (2) along each branch $Y_{i}$ is defined by $t$ and $s_{i}=1/t_{i}$. Due to the weighted blowup of the ideal $(x_{i},t^{a_{i}})$, where $a_{i}=\ell/(m_{i}+1)$, the restriction of $(t,s_{i})$ to $Y_{i}$ picks up the point $y_{i}$ defined by $s_{i}=0$ with coefficient $1/a_{i}=(m_{i}+1)/\ell$, and the last equivalence relation of divisor classes follows from $\sum_{i=1}^{n}(m_{i}+1)y_{i}\sim\omega_{Y}^{\log}$ in the construction of the contraction map $f$. Alternatively, we can compute the divisor at infinity by using [Pin77, Section 5, Formula ($\ast$)]. Indeed, in the family $\tilde{\mathcal{C}}$, we can further blow up each node of type $x_{i}t_{i}=t^{a_{i}}$ in the central fiber for $i=1,\ldots,n$, replacing it by a chain of rational curves of length $a_{i}$. The resulting family thus provides a minimal good resolution of $\mathcal{C}$ in the setting of [Pin77, Section 2], where the related invariants can be read off easily. Finally, note that $\omega_{Y}^{\log}$ is the restriction of the (relatively) ample line bundle $\mathcal{L}$ to $Y$, where $f^{*}\mathcal{L}\cong\tilde{\mathcal{L}}$ induces the contraction as in (4). Therefore, $\omega_{Y}^{\log}$ is ample. The same conclusion holds for $C$ and its divisor at infinity in the negatively graded miniversal deformation space of $(Y,q)$; see [Loo84b, (A.5)]. ∎
@@ 2.4. The boundary of $\mathbb{P}{\rm Def}^{-}(Y,q)$
As seen in Example 2.1, besides smooth elliptic curves in the miniversal deformation space of the ordinary cusp, the rational nodal curve also appears. In general, given a Gorenstein curve singularity $(Y,q)$ with $\mathbb{G}_{m}$-action, a natural question is to determine which singular curves can appear in $\mathbb{P}{\rm Def}^{-}(Y,q)$. This question has been studied in various special situations, mostly for unibranched monomial singularities with symmetric semigroups, for local complete intersections, and for low genus cases; see [Stö93, CS13, CF23, CM25, Del90, LP19, Bat22]. Utilizing the ideas in Theorem 2.6, we can provide a unified description to answer this question.
Analogous to the usual stability condition for nodal curves, we will use the following notion for Gorenstein curves. Let $(C,p_{1},\ldots,p_{n})$ be a reduced Gorenstein curve with $n$ marked points $p_{1},\ldots,p_{n}$ contained in the smooth locus of $C$. We say that $(C,p_{1},\ldots,p_{n})$ is a stable pointed Gorenstein curve if $\omega_{C}^{\log}=\omega_{C}(p_{1}+\cdots+p_{n})$ is an ample line bundle class. Furthermore, given a partition $\mu=(m_{1},\ldots,m_{n})$ of 2g22g-2, we say that the pointed curve $(C,p_{1},\ldots,p_{n})$ is of type $\mu$ if $\sum_{i=1}^{n}m_{i}p_{i}$ is a canonical divisor of $C$.
@@
Conversely, suppose $(C,p_{1},\ldots,p_{n})$ is a stable pointed Gorenstein curve with a differential $\omega$ as a section of the dualizing bundle of $C$, where the zeros of $\omega$ yield the canonical divisor $\sum_{i=1}^{n}m_{i}p_{i}$ of type $\mu$. Since each $p_{i}$ is a smooth point of $C$, the construction of the weighted blowup in Section 2.2 can be carried out in the same way. Additionally, $\omega_{C}^{\log}$ is ample by assumption, which implies that the twisted relative log dualizing bundle $\tilde{\mathcal{L}}$ can be defined as in (3) and remains relatively semiample. Indeed, let $\tilde{\eta}$ be the multi-scale differential associated to the limit of $\omega$ in the central fiber of the weighted blowup. Then $\tilde{\eta}$ as a relative section of $\tilde{\mathcal{L}}$ restricted to the central fiber has zeros only at the proper transforms $y_{i}$ of $p_{i}$. On the other hand, we can take a general section of the ample line bundle $\omega_{C}^{\log}$ such that it does not vanish at any $p_{i}$, and hence as a relative section of $\tilde{\mathcal{L}}$ it does not vanish at any $y_{i}$. Since $\tilde{\mathcal{L}}$ is a subbundle of $\omega_{\tilde{\pi}}^{\log}\otimes\mathcal{O}_{\tilde{\mathcal{C}}}(\ell(\tilde{\mathcal{C}}_{0}+\sum_{i=1}^{n}E_{i}))$ and the base of the family is a DVR, it follows that $\tilde{\pi}_{*}\tilde{\mathcal{L}}^{k}$ is locally free for $k\gg 0$. Therefore, the contraction in (4) holds.
We can also justify that the resulting singularity $q$ in the central fiber $Y$ after the contraction is Gorenstein. Indeed, the Residue Theorem holds for reduced singular curves as well; see [LP19, Lemma 1.1.2]. Additionally, by assumption $\tilde{\mathcal{C}}$ is Gorenstein in codimension $1$, and it also satisfies Serre’s condition $S_{2}$ since the base is smooth and the fibers are reduced curves. It follows that the dualizing sheaf of $\tilde{\mathcal{C}}$ can be identified with the sheaf associated to a canonical divisor; see [Kov13, Section 5]. Consequently, the same argument as in [CC26, Section 2] implies that the multi-scale differential $\tilde{\eta}$ as the limit of $\omega$ can descend to $Y$ and locally generate its dualizing sheaf at the singularity $q$, since the zeros of $\tilde{\eta}$ are along the divisor at infinity which are disjoint from $q$. The construction of the blowup and contraction as in Section 2.2 is $\mathbb{G}_{m}$-equivariant. Therefore, $(Y,q)$ is a Gorenstein curve with $\mathbb{G}_{m}$-action of type $\mu$, and its stability follows from Theorem 2.6 (v).
@@
Finally, the isomorphism of graded rings as in (7) follows from [Loo84b, (A.5)], where the class of the divisor at infinity is just $\omega^{\log}/\ell$. Indeed, the right-hand side of (5) can be rewritten as
$\ \left\{s\in H^{0}(C,m\omega_{C}^{\log})\mathrel{\Big|}(s)_{0}\geq\sum_{i=1}^{n}\frac{\lambda}{a_{i}}p_{i}\right\}$
$=$ $\ \left\{s\in H^{0}\left(C,m\omega_{C}^{\log}-\sum_{i=1}^{n}\frac{\lambda}{a_{i}}p_{i}\right)\right\}$
$=$ $\ H^{0}\left(C,(m\ell-\lambda)C_{\infty}\right),$
where $C_{\infty}=\sum_{i=1}^{n}(m_{i}+1)p_{i}/\ell=\sum_{i=1}^{n}p_{i}/a_{i}$ is the divisor at infinity for $C$. ∎
As an example, consider the special case when $(Y,q)$ is a monomial singularity whose gap sequence contains 2g12g-1; see Section 3.1 below for the related definitions. The following description of the boundary of $\mathbb{P}{\rm Def}^{-}(Y,q)$ was first established by Stöhr [Stö93] through a delicate study of Gröbner bases and syzygies for canonically embedded Gorenstein curves. Now using Theorem 2.7, we can give a concise and conceptual explanation as follows.
Corollary 2.8.
Let $(Y,q)$ be a monomial curve singularity of genus $g$ whose gap sequence $G$ contains 2g12g-1. Then the boundary of $\mathbb{P}{\rm Def}^{-}(Y,q)$ parameterizes stable pointed integral Gorenstein curves $(C,p)$ where the Weierstrass gap sequence of $p$ in $C$ is equal to $G$.
Proof.
In the case of a monomial singularity, the graded ring structure in (7) is determined by the corresponding semigroup that defines the singularity, whose complement is thus the gap sequence $G$; see Section 3.1 for more details. Note that $2g-1\in G$ is equivalent to that (2g2)p(2g-2)p is a canonical divisor, which corresponds to $\mu=(2g-2)$ in our notation. Therefore, the desired claim follows as a special case of Theorem 2.7. Finally, the irreducibility of $C$ is due to the ampleness of $\omega_{C}(p)$; otherwise if $C$ is reducible, then at least one of its irreducible components does not contain $p$, and hence $\omega_{C}(p)\sim(2g-1)p$ would be trivial restricted to that component, which contradicts that $\omega_{C}(p)$ is ample on $C$. ∎
@@ 3.1.4. Unibranched planar singularities
where $p$ and $q$ are coprime. Its $\mathbb{P}{\rm Def}^{-}_{s}$ corresponds to the locus in the stratum $\mathbb{P}\mathcal{H}(2g-2)$, where g=(p1)(q1)/2g=(p-1)(q-1)/2, with the gap sequence $G=\{b_{1},\ldots,b_{g}\}$ consisting of positive integers that cannot be expressed as $pi+qj$ with $i,j\geq 0$. In this case,